1. A 0.2500 g sample of a compound contain carbon, hydrogen and oxygen undergoes  complete combustion to produce 0.3664 g of CO2 and 0.1500 g of H2O. Determine the  empirical formula of this compound. [5 marks]


1
Expert's answer
2021-10-10T10:17:09-0400

From the question given above, the following data were obtained:


Mass of compound = 0.25 g


Mass of CO₂ = 0.3664 g


Mass of H₂O = 0.15 g


Empirical formula =?


Next, we shall determine the mass of carbon, hydrogen and oxygen present in the compound. This can be obtained as follow:


For Carbon (C):


Mass of CO₂ = 0.3664 g


Molar mass of CO₂ = 12 + (2×16) = 44 g/mol


Mass of C = 12/44 × 0.3664


Mass of C = 0.1



C= 0.1/12= 0.0083

H= 0.02/1= 0.02

O= 0.13/16= 0.0081

Divide by the smallest

C= 0.0083/0.0081= 1

H = 0.02/0.0081= 2.47

O= 0.0081/0.0081= 1

Multiply by 2 to express in whole number

C = 1× 2= 2

H = 2.47× 2= 5

O = 1× 2= 2

Empirical formula : C2H5O2